Week of 201005

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes

Monday

Attendance:

  • remote: Kate (chair, DB), Julia (WLCG), Maarten (WLCG, ALICE), Borja (monitoring), Marko (LHCb), Vincent (security), Andrew (TRIUMF), Andrew (NL-T1), Elena (CNAF), Xavier (KIT), Darren (RAL), Laurence (computing), Dave (FNAL)

Experiments round table:

Sites / Services round table:

  • ASGC: nc
  • BNL: nc
  • CNAF:
    • Scheduled Downtime for LHCb tomorrow to OS and GPFS upgrade. GOCDB link
    • Other scheduled downtimes:
      • 13/10 8:00 - 14:00 CEST to upgrade storm-cms + HSM (OS and GPFS)
      • 20/10 8:00 - 14:00 CEST to upgrade storm-atlas + HSM (OS and GPFS)
  • EGI: nc
  • FNAL: NTR
  • IN2P3: NTR
  • JINR: nc
  • KISTI: NTR
  • KIT:
    • Reminder for GridKa site downtime starting tomorrow 8 a.m. CEST.
    • Connectivity issues for ATLAS' transfers involving atlassrm-kit.gridka.de. In preparation for the upcoming downtime, we had installed new machines that will replace ATLAS' current dCache door nodes. Unfortunately dCache was running on those new hosts and consequently SRM tried to redirect clients to them, which failed. That was fixed early this morning.
  • NDGF: nc
  • NL-T1: NTR
  • NRC-KI: nc
  • OSG: nc
  • PIC: nc
  • RAL: NTR
  • TRIUMF: NTR

  • CERN computing services:
    • OTG:0059477 - CEs ce507.cern.ch & ce508.cern.ch upgrade to CentOS7 next Monday. Will be drained ahead of intervention. GOC entry has been added.
  • CERN storage services: nc
  • CERN databases: NTR
  • GGUS: NTR
  • Monitoring:
    • Draft availability reports for September sent around
  • MW Officer: nc
  • Networks: NTR
  • Security: NTR

AOB:

Edit | Attach | Watch | Print version | History: r16 < r15 < r14 < r13 < r12 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r16 - 2020-10-05 - KateDziedziniewicz
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback